OpenStudy (little_miss):

What is the simplified expression for 4 power 4 multiplied by 4 power 3 over 4 power 5?

OpenStudy (little_miss):

4^0 4^1 4^2 4^3

OpenStudy (steve816):

\[\huge 4^4*\frac{ 4^3 }{ 4^5 }\] Is this your question?

OpenStudy (allauhakbar):

D My school was cancer

OpenStudy (error1603):

D

OpenStudy (steve816):

IT IS NOT D!!!

OpenStudy (little_miss):

no its 4^4 x 4^3 4^5

OpenStudy (steve816):

Like what I have?

OpenStudy (little_miss):

well i guess sorry not good at math

OpenStudy (steve816):

Did I write your question correctly? So I can answer it.

OpenStudy (little_miss):

yes

OpenStudy (allauhakbar):

So it's D?

OpenStudy (little_miss):

I thought it was B

OpenStudy (little_miss):

No C

OpenStudy (steve816):

IT IS C

OpenStudy (steve816):

\[\huge 4^4 * 4^{-2} = 4^2\]

OpenStudy (little_miss):

YAY THANK YOU

OpenStudy (steve816):

No problem :)

OpenStudy (error1603):

I tried! I'm dumb!